3x^2+11x=4. I was told the answer was -3 2/3. I am stuck. I don't see how that was gotten.
Answer by stanbon(48558)

3x^2+11x=4
3x^2+11x-4=0
3x^2+12x-x-4=0
3x(x+4)-1(x+4)=0
(x+4)(3x-1)=0
x=-4 or x=1/3
Cheers,
Stan H.